OpenStudy (anonymous):

Khloe has a savings account that earns 3.13% compounded monthly. If she has $2,378 in the account today, what will the account balance be in 11 years? $2,447.13 $2,452.43 $3,337.70 $3,353.86

OpenStudy (anonymous):

@jim_thompson5910 @phi

jimthompson5910 (jim_thompson5910):

Use FV = PV*(1+r/n)^(n*t) where FV = unknown PV = 2378 r = 0.0313 n = 12 (compounded 12 times a year) t = 11 years

OpenStudy (phi):

any luck? Jim wrote it out (like setting the dinner table. Now serve the dinner), that is put in the numbers. You will want to use a calculator.

OpenStudy (anonymous):

Xavier deposits $6 daily into an interest-bearing account to save for renovations to his bathroom. The account earns 4.57% which compounds annually. What is the present value of the investment if Xavier renovates his bathroom in five years? $11,785.32 $10,033.83 $7,843.83 $9,595.32 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you familiar with annuities at all?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But I am not sure whether to use PVOA or PVAD

jimthompson5910 (jim_thompson5910):

PVOA = present value of ordinary annuity PVAD = present value of annuity due

jimthompson5910 (jim_thompson5910):

they're basically the same but the annuity due is where the payment is made at the beginning of the period (eg: rent)

OpenStudy (anonymous):

Yeah I know what they are but I am not which one to use in this scneario

jimthompson5910 (jim_thompson5910):

in this case, you use PVOA because the payment is made at the end of the period

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So I am trying to figure out how would I plug in the values

OpenStudy (anonymous):

Because Xavier is depositing 6 dollars daily

jimthompson5910 (jim_thompson5910):

hint: assume 365 days in a year (6 dollars per day) * (365 days per year) = 6*365 = 2190 dollars per year

jimthompson5910 (jim_thompson5910):

if the payment frequency doesn't match with the compounding frequency, that is your first goal: to get them to line up somehow

jimthompson5910 (jim_thompson5910):

so instead of thinking "6 dollars per day", think "2190 dollars per year" because the money is compounded annually

OpenStudy (anonymous):

would that be then used as the C in the equation

jimthompson5910 (jim_thompson5910):

yeah that's the cash flow C

OpenStudy (anonymous):

then the interest rate would be 0.0457/12 right?

OpenStudy (anonymous):

and 12 would be n right?

jimthompson5910 (jim_thompson5910):

the /12 part is only if you compound montly

jimthompson5910 (jim_thompson5910):

but we're actually compounding annually

OpenStudy (anonymous):

so????

jimthompson5910 (jim_thompson5910):

what formula are you using again? can you draw it out?

OpenStudy (anonymous):